OpenStudy (anonymous):

simplify 6r^2+6r/30r^2+36r

OpenStudy (anonymous):

\[6r ^{2}+6r/30r ^{2}+36r\] is this what you mean?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

I suck at polynomials, factoring and radicals

jimthompson5910 (jim_thompson5910):

\[\large \frac{6r^2+6r}{30r^2+36r}\] \[\large \frac{6r(r+1)}{6r(5r+6)}\] \[\large \frac{\cancel{6r}(r+1)}{\cancel{6r}(5r+6)}\] \[\large \frac{r+1}{5r+6}\] So \[\large \frac{6r^2+6r}{30r^2+36r}=\frac{r+1}{5r+6}\]

jimthompson5910 (jim_thompson5910):

In the second step, I factored out the GCF 6r and after that it divides and cancels out

OpenStudy (anonymous):

ok I see where you got the 5r+6 from